A car starts from a check point P and comes to rest at a check point Q,8km away in 4 minutes. It has first, a uniform acceleration for 50 seconds, then a constant speed, and brought to rest with uniform retardation after 30 seconds. Sketch a velocity time graph for the motion and find,maximum speed, acceleration and retardation.
1
Expert's answer
2019-10-22T10:25:02-0400


Maximum speed:


"v=40\frac{m}{s}"

Acceleration:


"a_1=\frac{v}{t_1}=\frac{40}{50}=0.8\frac{m}{s^2}"

"a_2=\frac{v}{t_2}=\frac{40}{30}=1.33\frac{m}{s^2}"